# Surveying optically addressable spin qubits for quantum information and sensing technology

### v2

@@ 2.3 Molecular FOAMs
Further organic systems have demonstrated potential as FOAMs, though, to our knowledge, pulsed optically detected experiments have yet to be performed. For example, the room temperature steady-state ODMR contrast of a 1 crystal of pentacene-doped picene has been measured at 15, representing a potentially significant improvement over the PTP matrix [130]. Steady-state ODMR signals have also previously been reported at 2 K for (perdeutero)tetracene, 1,2-benzathracene, 1,2,3,4-dibenzanthracene [131] and dinaphtho-(2’,3’:1,2);(2",3":6,7)-pyrene [132]. Interestingly, work by Corvaja, Pasimeni and Giometti et al., have shown that even highly spin-dense charge-transfer (CT) co-crystals can exhibit bright room-temperature ODMR signals. Co-crystals comprised of donors such as biphenyl, fluorene, phenazine and acceptors such as 7,7’:8,8’-tetracyanoquinodimethane (TCNQ) and 1,2,3,4-tetrafluoro-TCNQ (F4TCNQ) have been studied to elucidate their triplet state dynamics [133, 134]. These materials exhibit narrow resonance lines due to intermolecular site hopping of triplet excitons. The resonances for each site can become resolved at low temperatures where hopping is not thermodynamically favoured [133]. Their high spin densities (50) are highly advantageous for quantum sensing, where the a.c. sensitivity is proportional to the , though this is often concurrent with less robust quantum spin properties compared with dilute materials such as Pc:PTP. Moreover, these materials are also promising hosts for the study of exotic spin behaviours like singlet fission and triplet-triplet annihilation [135]. For example, at room-temperature, the triplet states of Phenazine:TCNQ are predominantly formed by singlet fission and using transient nutation ESR spectroscopy the authors estimate and times of 1 s and 600 ns, respectively, even at room temperature [136].
@@
2.3.2 d-block molecular spin centres
Molecular FOAMs with d-block elements have only recently been demonstrated with Cr4+. The Cr4+ spin-centre was realised with several tolyl-based ligand systems where modifications to the ligand structure and corresponding ligand field strength can be used to control optical excitation frequency and the ZFS of the Cr4+ S=1 ground state [142]. To reduce intermolecular dipole coupling, the spin-active moiety is diluted in a matrix with the S=0 isostructural tin analogue in crystals grown from hexane solutions [18]. Optical spin polarisation is then achieved by exciting molecular spins from the S=1 state to the S=0 manifold, which fluorescently decays within a few microseconds to favourably populate the T± states. Initial pulsed ODMR experiments were performed on Cr4+(o-tolyl)4 spins, which benefit from an E=0 triplet and a relatively long of 0.22 ms and of 640 ns at 5 K. This permits several optical cycles to build spin polarisation and enhance the optical contrast up to 14. The coherent properties can be significantly improved using a non-isostructural Sn(4-fluoro-2-methylphenyl)4 host matrix [143]. Here, the ZFS is significantly increased, giving rise to clock transitions similar to those realised by the so-called "zero first-order Zeeman" (ZEFOZ) method, initially pioneered for lanthanide spin systems (see below). As a result, and are increased to 1.210.02 ms and 10.60.2 s at 5 K, respectively, while the pulsed contrast can be improved up to 65.
@@
2.3.3 f-block molecular spin centres
@@
In the solid-state, Y2SiO5 has been favoured as host matrix due to the ability to grow large crystals with excellent optical properties using the Czochralski method [160] (Figure 9b). However, as the only naturally occurring isotope, 89Y harbours an I=1/2 nuclear spin that ultimately limits decoherence times. To reduce the impact of the spin bath, various decoupling techniques have emerged [161, 162, 45]. Perhaps the most successful is the so-called "ZEFOZ" technique, whereby a magnetic field is applied such that the magnetic-field dependence of the spin-transition frequency is very close to zero [163] (Figure 9d). In this "clock transition" regime, spins are first-order insensitive to small fluctuations in local magnetic fields. Using this technique with Eu3+:Y2SiO5, Zhong et al., demonstrated it is possible to acquire >>100 ms at 2 K, where spin-phonon coupling is negligible. Remarkably, combined with dynamic decoupling methods was measured to be 37060 mins, reaching a critical milestone whereby the distance-dependent decoherence becomes less for spin-transport than it is during light-transport of quantum states [164].
Interestingly, the larger magnetic moment of Pr3+ can give rise to "frozen core"-type behaviour, whereby local Y-spins become dephased from the bulk crystal and hence exhibit reduced dephasing influence on the Pr3+ spins [161]. Equall et al., measured homogenous field-dependent and crystal structure site-dependent linewidths between 2.5 and 0.85 kHz, and a corresponding as high as 377 s at 1.4 K [50]. Combined with the ZEFOZ method, the can reach 82 ms at 1.5 K [161] (see example Figure 9e) and with further dynamical decoupling up to 1 min can be achieved, approaching the population lifetime limit [165]. Coherence times can also be enhanced using a host matrix whereby the principal host ion (e.g., Y) has a more closely matched ionic radius to the dopant, leading to reduced crystallographic distortions. For example, in a Pr3+:La2(WO4)3 system a of 1587 ms has been measured at 4 K [51].
@@
171Yb3+ holds a unique position amongst the lanthanide ions discussed so far due to its S=1/2 and I=1/2 electron spin and hyperfine structure, resulting in a simple 4-level system. Of the 171Yb3+-doped materials [53], 171Yb3+:Y2SiO5 appears to exhibit the longest spin coherence times. This material was initially studied by X-band EPR spectroscopy and presented with an electron of 5 seconds at 2.5 K with an applied field of 90 mT. quickly increases above 4 K due to Raman relaxation where becomes ultimately limited by [52]. At 2.5 K, was optimised at 1 T to 73 s and improved further by dynamical decoupling reaching 550 s. In the same experiments, the authors record nuclear and at 4.5 K of 4 and 0.35 ms, respectively. Using an optical approach, Ortu et al., improved the coherent properties of 171Yb3+:Y2SiO5 by employing the ZEFOZ method such that the electron remains above 100 s at 5.6 K and the nuclear extends to 1 ms [158]. Using a different approach, Welinski et al., demonstrated that coherence can also be extended by first polarising host nuclear spins through spin diffusion. At 2 K, the authors first excite 171Yb spins before allowing them to equilibrate over a few seconds through spectral diffusion over the inhomogeneous linewidth. The result is an effective "hole burning" in the absorption spectrum of 171Yb3+:Y2SiO5 and up to 90 nuclear spin polarisation, thereby effectively generating mK spin temperatures and improving the optical from 0.3 to 0.8 ms [168].
3 Optimising Performance of Spin-Qubit Materials
@@
From our review, it is clear that spin-based qubit candidates demonstrate potential in several fields of quantum technology. However, unsurprisingly, their spin coherence lifetimes are limited by temperature and the spin-bath dependence of the spin properties and optical contrast between different spin states. Across all material platforms, there is significant magnetic inhomogeneity that emerges from random local spin environments. To understand the extent to which inhomogeneity infects different materials, it could be instructive to consider the ratio of / (Figure 10, the so-called "inhomogeneity parameter"). Hence, a low ratio indicates that is close to and the materials are limited by dynamic, not static, sources of decoherence. Using the available data where and were measured under similar conditions, there appears to emerge a distinct advantage for molecules and van der Waals materials at higher temperatures despite their lack of isotopic enrichment. This likely stems from the use of molecular crystals and the inherent 2D-order associated materials such as hBN, which helps to helps to ensure that all molecules "feel" the same magnetic environment. Interestingly, despite the significant difficulties associated with synthesising aligned 3D-spin defects in diamond, SnV-diamond (measured as single spins) exhibits the lowest ratio at low temperatures. Considering its robust quantum spin parameters, SnV-diamond could be a leading candidate for low-temperature applications.

### v3

@@ II.3 Molecular FOAMs
Further organic systems have demonstrated potential as FOAMs, though, to our knowledge, pulsed optically detected experiments have yet to be performed. For example, the room temperature steady-state ODMR contrast of a 1 crystal of pentacene-doped picene has been measured at 15, representing a potentially significant improvement over the PTP matrix [130]. Steady-state ODMR signals have also previously been reported at 2 K for (perdeutero)tetracene, 1,2-benzathracene, 1,2,3,4-dibenzanthracene [131] and dinaphtho-(2’,3’:1,2);(2”,3”:6,7)-pyrene [132]. Interestingly, work by Corvaja, Pasimeni and Giometti et al., have shown that even highly spin-dense charge-transfer (CT) co-crystals can exhibit bright room-temperature ODMR signals. Co-crystals comprised of donors such as biphenyl, fluorene, phenazine and acceptors such as 7,7’:8,8’-tetracyanoquinodimethane (TCNQ) and 1,2,3,4-tetrafluoro-TCNQ (F4TCNQ) have been studied to elucidate their triplet state dynamics [133, 134]. These materials exhibit narrow resonance lines due to intermolecular site hopping of triplet excitons. The resonances for each site can become resolved at low temperatures where hopping is not thermodynamically favoured [133]. Their high spin densities (50) are highly advantageous for quantum sensing, where the a.c. sensitivity is proportional to the , though this is often concurrent with less robust quantum spin properties compared with dilute materials such as Pc:PTP. Moreover, these materials are also promising hosts for the study of exotic spin behaviours like singlet fission and triplet-triplet annihilation [135]. For example, at room-temperature, the triplet states of Phenazine:TCNQ are predominantly formed by singlet fission and using transient nutation ESR spectroscopy the authors estimate and times of 1 s and 600 ns, respectively, even at room temperature [136].
@@
II.3.2 d-block molecular spin centres
Molecular FOAMs with d-block elements have only recently been demonstrated with Cr4+. The Cr4+ spin-centre was realised with several tolyl-based ligand systems where modifications to the ligand structure and corresponding ligand field strength can be used to control optical excitation frequency and the ZFS of the Cr4+ S=1 ground state [142]. To reduce intermolecular dipole coupling, the spin-active moiety is diluted in a matrix with the S=0 isostructural tin analogue in crystals grown from hexane solutions [18]. Optical spin polarisation is then achieved by exciting molecular spins from the S=1 state to the S=0 manifold, which fluorescently decays within a few microseconds to favourably populate the T± states. Initial pulsed ODMR experiments were performed on Cr4+(o-tolyl)4 spins, which benefit from an E=0 triplet and a relatively long of 0.22 ms and of 640 ns at 5 K. This permits several optical cycles to build spin polarisation and enhance the optical contrast up to 14. The coherent properties can be significantly improved using a non-isostructural Sn(4-fluoro-2-methylphenyl)4 host matrix [143]. Here, the ZFS is significantly increased, giving rise to clock transitions similar to those realised by the so-called ”zero first-order Zeeman” (ZEFOZ) method, initially pioneered for lanthanide spin systems (see below). As a result, and are increased to 1.210.02 ms and 10.60.2 s at 5 K, respectively, while the pulsed contrast can be improved up to 65.
@@
II.3.3 f-block molecular spin centres
@@
In the solid-state, Y2SiO5 has been favoured as host matrix due to the ability to grow large crystals with excellent optical properties using the Czochralski method [160] (Figure 9b). However, as the only naturally occurring isotope, 89Y harbours an I=1/2 nuclear spin that ultimately limits decoherence times. To reduce the impact of the spin bath, various decoupling techniques have emerged [161, 162, 45]. Perhaps the most successful is the so-called ”ZEFOZ” technique, whereby a magnetic field is applied such that the magnetic-field dependence of the spin-transition frequency is very close to zero [163] (Figure 9d). In this ”clock transition” regime, spins are first-order insensitive to small fluctuations in local magnetic fields. Using this technique with Eu3+:Y2SiO5, Zhong et al., demonstrated it is possible to acquire ¿¿100 ms at 2 K, where spin-phonon coupling is negligible. Remarkably, combined with dynamic decoupling methods, was measured to be 37060 mins, reaching a critical milestone whereby the distance-dependent decoherence becomes less for spin-transport than it is during light-transport of quantum states [164].
Interestingly, the larger magnetic moment of Pr3+ can give rise to ”frozen core”-type behaviour, whereby local Y-spins become dephased from the bulk crystal and hence exhibit reduced dephasing influence on the Pr3+ spins [161]. Equall et al., measured homogenous field-dependent and crystal structure site-dependent linewidths between 2.5 and 0.85 kHz, and a corresponding as high as 377 s at 1.4 K [50]. Combined with the ZEFOZ method, the can reach 82 ms at 1.5 K [161] (see example Figure 9e) and with further dynamical decoupling up to 1 min can be achieved, approaching the population lifetime limit [165]. Coherence times can also be enhanced using a host matrix whereby the principal host ion (e.g., Y) has a more closely matched ionic radius to the dopant, leading to reduced crystallographic distortions. For example, in a Pr3+:La2(WO4)3 system a of 1587 ms has been measured at 4 K [51].
@@
171Yb3+ holds a unique position amongst the lanthanide ions discussed so far due to its S=1/2 and I=1/2 electron spin and hyperfine structure, resulting in a simple 4-level system. Of the 171Yb3+-doped materials [53], 171Yb3+:Y2SiO5 appears to exhibit the longest spin coherence times. This material was initially studied by X-band EPR spectroscopy and presented with an electron of 5 seconds at 2.5 K with an applied field of 90 mT. quickly increases above 4 K due to Raman relaxation where becomes ultimately limited by [52]. At 2.5 K, was optimised at 1 T to 73 s and improved further by dynamical decoupling reaching 550 s. In the same experiments, the authors record nuclear and at 4.5 K of 4 and 0.35 ms, respectively. Using an optical approach, Ortu et al., improved the coherent properties of 171Yb3+:Y2SiO5 by employing the ZEFOZ method such that the electron remains above 100 s at 5.6 K and the nuclear extends to 1 ms [158]. Using a different approach, Welinski et al., demonstrated that coherence can also be extended by first polarising host nuclear spins through spin diffusion. At 2 K, the authors first excite 171Yb spins before allowing them to equilibrate over a few seconds through spectral diffusion over the inhomogeneous linewidth. The result is an effective ”hole burning” in the absorption spectrum of 171Yb3+:Y2SiO5 and up to 90 nuclear spin polarisation, thereby effectively generating mK spin temperatures and improving the optical from 0.3 to 0.8 ms [168].
III Optimising Performance of Spin-Qubit Materials
@@
From our review, it is clear that spin-based qubit candidates demonstrate potential in several fields of quantum technology. However, unsurprisingly, their spin coherence lifetimes are limited by temperature and the spin-bath dependence of the spin properties and optical contrast between different spin states. Across all material platforms, there is significant magnetic inhomogeneity that emerges from random local spin environments. To understand the extent to which inhomogeneity infects different materials, it could be instructive to consider the ratio of / (Figure 10, the so-called ”inhomogeneity parameter”). Hence, a low ratio indicates that is close to and the materials are limited by dynamic, not static, sources of decoherence. Using the available data where and were measured under similar conditions, there appears to emerge a distinct advantage for molecules and van der Waals materials at higher temperatures despite their lack of isotopic enrichment. This likely stems from the use of molecular crystals and the inherent 2D-order associated materials such as hBN, which helps to ensure that all molecules ”feel” the same magnetic environment. Interestingly, despite the significant difficulties associated with synthesising aligned 3D-spin defects in diamond, SnV-diamond (measured as single spins) exhibits the lowest ratio at low temperatures. Considering its robust quantum spin parameters, SnV-diamond could be a leading candidate for low-temperature applications.